Select the lightest W shape to support a uniformly distributed load of 80 kN/m and concentrated loads of 26kN located 4.5 m from the end of the beam. The beam is 16 m, Deflection is not to exceed span/280.
Answer #1

Draw FBD of beam 12 Belen 26kw V w Solm mmmmmmmm lec k 40 sm sm Ks = 64-512 Ka = 5s-675 16 m The vou get replaced by roint loNow we deplop to deflecrion eq? M2x72 658 05975 <*> - 26 {æ-4.57- 80 <u>2 04221cm El d2y = Mau dx2 Ez dyldne = 658-6575 <2}}-we need to define first maximum strope deflection at maximum deflection slope will be zero let it occurs b/w cand (3) El dupldel cellow) = -4260 (Shes = 16 (_m) T allow = -0.05714m] 280 Ymax = (d allow = – 69740.262 =-0.05714 m E ? Es 200 x 105 od le